Question 6 18 pts In circuit 1, four resistors are connected purely in parallel to a battery. Three resistors are unknown and one of the resistors is 6 Ohms. In circuit 2, the resistor whose resistance is known (6 Ohms is removed from the circuit and the other three resistors are connected purely in parallel to the same battery. The power output of the battery in circuit 1 is 5.1 Watts more than the power output of the battery in circuit 2. What is the voltage of the battery in Volts?
Answer #1

for circuit 1 :

P1 = V2/R1 + V2/R2 + V2/R3 + V2/6

for circuit 2

P2 = V2/R1 + V2/R2 + V2/R3 = V2 (1/R1 + 1/R2 + 1/R3) = V2 /Rp =

given that :

P1 = 5.1 + P2

V2/R1 + V2/R2 + V2/R3 + V2/6 = 5.1 + V2/R1 + V2/R2 + V2/R3

V2/6 = 5.1

V = sqrt(6 x 5.1) = 5.53 volts